# Magic Boundaries of 3D Color Codes

### v3

@@ I Introduction
Among a large family of topological stabilizer codes, the color code has obtained significant interest due to its capability for fault-tolerant quantum operations, particularly transversal logical gates [4]. Logical gates from the Clifford group can be implemented transversally in the 2D color code by applying a unitary operation to each physical qubit. Additionally, it is worth noting that the 2D color code boundary can be used to distill magic states through code deformation [9], which can be utilized for universal quantum computing [10]. More generally, a $D$-dimensional color code admits transversal logical gates in the $D^{\text{th}}$ level of the Clifford hierarchy [11]. Furthermore, an equivalence is established between $D$-dimensional color codes and $D$ copies of $D$-dimensional toric codes [11]. From the TQFT perspective, a $D$-dimensional toric code corresponds to a ($D$+1)-dimensional $\mathbb{Z}_{2}$ gauge theory, while a $D$-dimensional color code corresponds to a ($D$+1)-dimensional $\mathbb{Z}_{2}^{D}$ gauge theory.
@@
I.1 Main results and ideas
@@
I.2 Outline
The paper is organized as follows. In Section II.1 and II.2, we review the definition of the 3D toric code, the 3D color code, and their corresponding topological excitations. In Section II.3, we review the definition of the Lagrangian subgroup, and In Section II.4, we introduce the unfolding unitary transformation that maps between the toric code and the color code without boundaries. In Section III, we introduce the unfolding unitaries on the systems with boundaries, and further define the $X$-boundary, the $Z$-boundary and the other boundaries with a mixture of $X$ and $Z$ types condensations. In Section IV, we construct the magic boundary from the 3D color code with $X$-boundaries, and show the exotic properties of this boundary explicitly. We further discuss the nested boundaries, which are boundaries with codimension-2 (1D) $S$-domain wall defects. In Section V, we draw the conclusion and have a discussion of our work, followed by an outlook towards future directions. A detailed description of the unfolding unitary transformation is discussed in Appx. A. Gauging the 3D Ising model in the presence of boundaries is discussed in Appx. B. A minimal model that has two $Z$-boundaries and four $X$-boundaries is constructed explicitly in Appx. C.
II Backgrounds
@@ II.1 The 3D toric code
In this paper we consider two types of elementary boundaries of the 3D toric code: the rough (e) boundary and the smooth (m) boundary. $Z$-strings can terminate on the rough boundary without generating any electric excitations at the intersections, while $X$-membranes can terminate on the smooth boundary without generating any flux loops at the intersections. It is equivalent to say that the $e$-particle can condense on the rough boundary, while the $m$-flux loop can condense on the smooth boundary. From a mathematical perspective, these condensations form the Lagrangian subgroup on the boundary, as detailed in [12]. We have a brief review of the Lagrangian subgroup in Sec. II.3.
@@
II.2 The 3D color code
The $3$-dimensional topological color code is defined on a lattice $\mathcal{L}(\mathcal{V},\mathcal{E},\mathcal{F},\mathcal{C})$ with the following constraints. First, each vertex on the lattice belongs to four different edges. Second, all the cells are four-colorable, which means they can be represented by four colors, and each pair of adjacent cells has different colors.
@@ II.2 The 3D color code
In general, one can choose arbitrary 4-colorable and 4-valent cellulations for the $3$D color code. For the sake of convenience, we select a particular cellulation for the three-dimensional topological color code, as depicted in Fig. 1. It’s important to note that the results we obtain in this paper apply generally and are not contingent on the specific cellulation chosen.
In the lattice shown in Fig. 1, we use yellow, green, purple, and red colors, $\{y,g,p,r\}$, to label the color of cells and the edges that connect cells with the same color, and we use two colors to label the plaquettes between two cells with different colors, namely $\{rp,rg,ry,pg,py,yg\}$.
@@
To avoid confusion, we use a single color to label these excitations. To create a single type of flux loop, one can apply $X$-operators on the corresponding membrane and the flux loop is located on the boundary of that membrane. For example, the $yg_{\mathbf{x}}$ flux loop is created by the following membrane operator.
@@
A more detailed discussion of the $X$-membrane operators and the flux loops can be found in Sec. III.2.
@@
in which $\mathbf{c}_{i},\mathbf{c}_{j},\mathbf{c}_{k}\in\mathbf{C}$.
@@
II.3 Gapped boundaries of topological phases
In the previous subsections, we learned that in the presence of boundaries, some excitations from the bulk of topological phases can condense on these boundaries, while others remain gapped. From the perspective of operators, this is equivalent to saying that the boundaries of certain excitation creation operators can terminate on such boundaries, resulting in no excitations being created at the intersections. The set of excitations that can condense on the boundary forms a group, referred to as the Lagrangian subgroup in finite abelian theories, as discussed in [22, 12].
@@ II.3 Gapped boundaries of topological phases
The set $\mathsf{M}$ that satisfies the above conditions are call the Lagrangian subgroup of $\mathsf{A}$.
@@
II.4 Unfolding the 3D Color Code without Boundaries
@@
Consider the lattice $\mathcal{L}$ associated with a $3$-dimensional topological color code, as we defined in Sec. II. The unfolding process of color $\mathbf{c}_{i},\mathbf{c}_{j},\mathbf{c}_{k}\in\mathbf{C}$, $i\neq j\neq k$ on the lattice and the map between stabilizers can be described as follows. First, apply local unitary operators to map all the Pauli operators on vertices to Pauli operator on the $\mathbf{c}_{i}$-, $\mathbf{c}_{j}$-, and $\mathbf{c}_{k}$-colored edges. Second, shrink all the $\mathbf{c}_{i}$-, $\mathbf{c}_{j}$-, and $\mathbf{c}_{k}$-colored cells sequentially to points. After this step, we obtain three copies of lattices with different colors. Third, disentangle and discard the qubits on vertices. The outcome of the above process yields three copies of 3D toric codes model defined on the $\mathbf{c}_{i}$-, $\mathbf{c}_{j}$-, and $\mathbf{c}_{k}$-colored lattices, respectively. In a broader sense, for a $D$-dimensional color code, selecting and shrinking $D$ distinct colors can lead to $D$ copies of the $D$-dimensional toric codes.
@@
  • By shrinking the green (yellow) cells, the lattice transforms into the following configuration. The “…” in the figure indicates the lattice is extended along the horizontal and vertical directions. The $X$-type stabilizers are located at vertices, and each $X$-stabilizer consists of the tensor product of $X$-operators on the nearest edges. The $Z$-type stabilizers are given by the tensor product of $Z$-operators along the edges of faces. [^2] Acute readers may notice that based on the shrinking rule, there are four qubits on each edge of the green (yellow) lattice. However, we can further apply local unitaries to decouple these degrees of freedom and leave only one qubit per edge, which are those of the 3D toric code. Details of this process is discussed in Appx. A
@@
An example of the transformations of stabilizers and lattice deformation on the purple lattice is summarized in Fig. 2, and a detailed description of the unfolding unitaries is discussed in Appx. A.
@@
III Unfolding the 3D color code with boundaries
Figure 3: An illustration of the equivalence between the 3D color code with different types of boundaries and three copies of toric codes with different alignments. (a) A 3D color code with color boundaries (each pair of boundaries can condense a distinct type of ($\mathbf{c}_{i})_{\mathbf{z}}$-particles.) is equivalent to three copies of toric codes with distinct alignments of boundaries. Therefore, each color boundary corresponds to a $(e_{i},m_{j},m_{k})$-boundary of three copies of 3D toric codes. (b) A 3D color code with two $Z$-boundaries and four $X$-boundaries is equivalent to three copies of toric codes with the same alignment of boundaries. Therefore, the $Z$-boundaries correspond to the $(e_{1},e_{2},e_{3})$-boundaries and the $X$-boundaries correspond to the $(m_{1},m_{2},m_{3})$-boundaries. A miminal model that can realize this boundary condition is explicitly constructed in Appx. C
@@ III Unfolding the 3D color code with boundaries
For the convenience of our discussion, we define $\mathbf{c}_{1}=y$, $\mathbf{c}_{2}=g$ and $\mathbf{c}_{3}=p$. In the remainder of this section, we will demonstrate, following the original proof in Ref. [11], that the $(e_{1},e_{2},e_{3})$ (all-rough) boundary of three copies of toric codes is equivalent to the $\{y_{\mathbf{z}},g_{\mathbf{z}},p_{\mathbf{z}}\}$-boundary ($Z$-boundary) of the 3D color code, and that the $(m_{1},m_{2},m_{3})$ (all-smooth) boundary is equivalent to the $\{pg_{\mathbf{x}},py_{\mathbf{x}},yg_{\mathbf{x}}\}$-boundary ($X$-boundary), up to local unitaries and entangling/disentangling ancilla qubits. We illustrate the equivalence between the 3D color code with two $Z$-boundaries and four $X$-boundaries, and three copies of the 3D toric codes with the same boundary alignment in Fig. 3(b). A minimal model capable of realizing this boundary condition is explicitly constructed in Appx. C. Moreover, we explicitly construct the boundary Hamiltonians for the other boundaries that can be classified by the Lagrangian subgroup.
III.1 The $Z$-boundary
@@
Our next step is to show that this boundary, up to local unitary transformations and entangling/disentangling ancilla qubits, is equivalent to the $(e_{1},e_{2},e_{3})$-boundary of three copies of 3D toric codes. We apply the unfolding unitaries introduced in Sec. II.4. To simplify the presentation of the boundary, we have opted for a different alignments from the one shown for the string operators. In this case, we consider the front and rear sides of the green (and purple) lattice to be the $Z$-boundaries, as opposed to the top and bottom sides as previously shown. The green sublattice is given by
@@ III.2 The $X$-Boundary
Similar to the case with $Z$-boundaries, three independent types of membrane operators can terminate on this boundary and commute with every stabilizer. However, the membranes of the 3D color code are not flat, and their shapes vary depending on the specific lattice. In our example, we display the configurations of these (logical) membrane operators in Fig. 4. We define the $X$-membrane operator on the $\mathbf{c}_{i}\mathbf{c}_{j}$-colored membrane as $\overline{X}_{\mathbf{c}_{i}\mathbf{c}_{j}}$ if it is also a logical operator, and define it as $\widetilde{X}_{\mathbf{c}_{i}\mathbf{c}_{j}}$ if it is not.
@@ III.3 Other boundaries
III.3.1 The $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$-boundaries
@@
III.3.2 The $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$-boundaries
@@
III.3.3 The folded boundaries
[…]

### v4

@@ 1 Introduction
Among a large family of topological stabilizer codes, the color code has obtained significant interest due to its capability for fault-tolerant quantum operations, particularly transversal logical gates [4]. Logical gates from the Clifford group can be implemented transversally in the 2D color code by applying a unitary operation to each physical qubit. Additionally, it is worth noting that the 2D color code boundary can be used to distill magic states through code deformation [9], which can be utilized for universal quantum computing [10]. More generally, a $D$-dimensional color code admits transversal logical gates in the $D^{\text{th}}$ level of the Clifford hierarchy [11]. Furthermore, an equivalence is established between $D$-dimensional color codes and $D$ copies of $D$-dimensional toric codes [11]. From the TQFT perspective, a $D$-dimensional toric code corresponds to a (DD+1)-dimensional $\mathbb{Z}_{2}$ gauge theory, while a $D$-dimensional color code corresponds to a (DD+1)-dimensional $\mathbb{Z}_{2}^{D}$ gauge theory.
@@
1.1 Main results and ideas
@@
1.2 Outline
The paper is organized as follows. In Section 2.1 and 2.2, we review the definition of the 3D toric code, the 3D color code, and their corresponding topological excitations. In Section 2.3, we review the definition of the Lagrangian subgroup, and In Section 2.4, we introduce the unfolding unitary transformation that maps between the toric code and the color code without boundaries. In Section 3, we introduce the unfolding unitaries on the systems with boundaries, and further define the $X$-boundary, the $Z$-boundary and the other boundaries with a mixture of $X$ and $Z$ types condensations. In Section 4, we construct the magic boundary from the 3D color code with $X$-boundaries, and show the exotic properties of this boundary explicitly in Subsection 4.5. We further discuss the nested boundaries in Subsection 4.7, which are boundaries with codimension-2 (1D) $S$-domain wall defects. In Section 5, we draw the conclusion and have a discussion of our work, followed by an outlook towards future directions. A detailed description of the unfolding unitary transformation is discussed in Appendix A. Gauging the 3D Ising model in the presence of boundaries is discussed in Appendix B. A minimal model that has two $Z$-boundaries and four $X$-boundaries is constructed explicitly in Appendix C.
2 Backgrounds
@@ 2.1 The 3D toric code
In this paper we consider two types of elementary boundaries of the 3D toric code: the rough (e) boundary and the smooth (m) boundary. $Z$-strings can terminate on the rough boundary without generating any electric excitations at the intersections, while $X$-membranes can terminate on the smooth boundary without generating any flux loops at the intersections. It is equivalent to say that the $e$-particle can condense on the rough boundary, while the $m$-flux loop can condense on the smooth boundary. From a mathematical perspective, these condensations form the Lagrangian subgroup on the boundary, as detailed in [12]. We have a brief review of the Lagrangian subgroup in Section 2.3.
@@
2.2 The 3D color code
The $3$-dimensional topological color code is defined on a lattice $\mathcal{L}(\mathcal{V},\mathcal{E},\mathcal{F},\mathcal{C})$ with the following constraints. First, each vertex on the lattice belongs to four different edges. Second, each cells can be assigned one of four colors such that each pair of adjacent cells has different colors.
@@ 2.2 The 3D color code
In general, one can choose arbitrary 4-colorable and 4-valent cellulations for the 33D color code. For the sake of convenience, we select a particular cellulation for the three-dimensional topological color code, as depicted in Fig. 1. It’s important to note that the results we obtain in this paper apply generally and are not contingent on the specific cellulation chosen.
In the lattice shown in Fig. 1, we use yellow, green, purple, and red colors, $\mathbf{c}_{i}\in\{y,g,p,r\}=\mathbf{C}$, to label the color of cells and the edges that connect cells with the same color, and we use two colors to label the plaquettes between two cells with different colors, namely $\{rp,rg,ry,pg,py,yg\}$.
@@
To avoid confusion, we use a single color to label these excitations. To create a single type of flux loop, one can apply $X$-operators on the corresponding membrane and the flux loop is located on the boundary of that membrane. This membrane is formed by plaquettes having two particular colors on the edges. For example, the $yg_{\mathbf{x}}$ flux loop is created by the following membrane operator, which is supported on plaquettes bounded by yellow and green edges.
@@
A more detailed discussion of the $X$-membrane operators and the flux loops can be found in Section 3.2.
@@
in which $\mathbf{c}_{i},\mathbf{c}_{j},\mathbf{c}_{k}\in\mathbf{C}$, where $\mathbf{c}_{i},\mathbf{c}_{j},\mathbf{c}_{k}$ denote color labels ranging over the set $\mathbf{C}$.
@@
2.3 Gapped boundaries of topological phases
In the previous subsections, we learned that in the presence of boundaries, some excitations from the bulk of topological phases can condense on these boundaries, while others remain gapped. From the perspective of operators, this is equivalent to saying that the boundaries of certain excitation creation operators can terminate on such boundaries, resulting in no excitations being created at the intersections where the membrane meets the boundary. The set of excitations that can condense on the boundary forms a group, referred to as the Lagrangian subgroup in finite abelian theories, as discussed in [22, 12].
@@ 2.3 Gapped boundaries of topological phases
The set $\mathsf{M}$ that satisfies the above conditions are call a Lagrangian subgroup of $\mathsf{A}$.
@@
2.4 Unfolding the 3D Color Code without Boundaries
@@
Consider the lattice $\mathcal{L}$ associated with a $3$-dimensional topological color code, as we defined in Section 2. The unfolding process of color $\mathbf{c}_{i},\mathbf{c}_{j},\mathbf{c}_{k}\in\mathbf{C}$, $i\neq j\neq k$ on the lattice and the map between stabilizers can be briefly summarized as follows. First, prepare one ancilla qubit on each edge colored $\mathbf{c}_{i}$, $\mathbf{c}_{j}$, or $\mathbf{c}_{k}$, where $\mathbf{c}_{i},\mathbf{c}_{j},\mathbf{c}_{k}\in\mathbf{C}$. Second, apply a local unitary operator $U=\bigotimes_{c}U_{c}$ on every cell with color $\mathbf{c}_{i},\mathbf{c}_{j},\mathbf{c}_{k}\in\mathbf{C}$ such that the original stabilizers in the 3D color code are mapped to three sets of stabilizers, each supported on one of the three corresponding sets of qubits on edges with different colors. Third, shrink all the $\mathbf{c}_{i}$-, $\mathbf{c}_{j}$-, and $\mathbf{c}_{k}$-colored cells sequentially to points. After this step, we obtain three copies of lattices with different colors. Fourth, disentangle and discard the qubits on vertices. The outcome of the above process yields three copies of 3D toric codes model defined on the $\mathbf{c}_{i}$-, $\mathbf{c}_{j}$-, and $\mathbf{c}_{k}$-colored lattices, respectively. In a broader sense, for a $D$-dimensional color code, selecting and shrinking $D$ distinct colors can lead to $D$ copies of the $D$-dimensional toric codes.
@@
  • By shrinking the green (yellow) cells, the lattice transforms into the following configuration. The “…” in the figure indicates the lattice is extended along the horizontal and vertical directions. The $X$-type stabilizers are located at vertices, and each $X$-stabilizer consists of the tensor product of $X$-operators on the nearest edges. The $Z$-type stabilizers are given by the tensor product of $Z$-operators along the edges of faces. [^2] Acute readers may notice that based on the shrinking rule, there are four qubits on each edge of the green (yellow) lattice. However, we can further apply local unitaries to decouple these degrees of freedom and leave only one qubit per edge, which are those of the 3D toric code. Details of this process is discussed in Appendix A
@@
An example of the transformations of stabilizers and lattice deformation on the purple lattice is summarized in Fig. 2, and a detailed description of the unfolding unitaries is discussed in Appendix A.
@@
3 Unfolding the 3D color code with boundaries
Figure 3: An illustration of the equivalence between the 3D color code with different types of boundaries and three copies of toric codes with different alignments. (a) A 3D color code with color boundaries (each pair of boundaries can condense a distinct type of ($\mathbf{c}_{i})_{\mathbf{z}}$-particles.) is equivalent to three copies of toric codes with distinct alignments of boundaries. Therefore, each color boundary corresponds to a $(e_{i},m_{j},m_{k})$-boundary of three copies of 3D toric codes. (b) A 3D color code with two $Z$-boundaries and four $X$-boundaries is equivalent to three copies of toric codes with the same alignment of boundaries. Therefore, the $Z$-boundaries correspond to the $(e_{1},e_{2},e_{3})$-boundaries and the $X$-boundaries correspond to the $(m_{1},m_{2},m_{3})$-boundaries. A miminal model that can realize this boundary condition is explicitly constructed in Appendix C
@@ 3 Unfolding the 3D color code with boundaries
For the convenience of our discussion, we define $\mathbf{c}_{1}=y$, $\mathbf{c}_{2}=g$ and $\mathbf{c}_{3}=p$. In the remainder of this section, we will demonstrate, following the original proof in Ref. [11], that the $(e_{1},e_{2},e_{3})$ (all-rough) boundary of three copies of toric codes is equivalent to the $\{y_{\mathbf{z}},g_{\mathbf{z}},p_{\mathbf{z}}\}$-boundary ($Z$-boundary) of the 3D color code, and that the $(m_{1},m_{2},m_{3})$ (all-smooth) boundary is equivalent to the $\{pg_{\mathbf{x}},py_{\mathbf{x}},yg_{\mathbf{x}}\}$-boundary ($X$-boundary), up to local unitaries and entangling/disentangling ancilla qubits. We illustrate the equivalence between the 3D color code with two $Z$-boundaries and four $X$-boundaries, and three copies of the 3D toric codes with the same boundary alignment in Fig. 3(b). A minimal model capable of realizing this boundary condition is explicitly constructed in Appendix C. Moreover, we explicitly construct the boundary Hamiltonians for the other boundaries that can be classified by the Lagrangian subgroup.
3.1 The $Z$-boundary
@@
Our next step is to show that this boundary, up to local unitary transformations and entangling/disentangling ancilla qubits, is equivalent to the $(e_{1},e_{2},e_{3})$-boundary of three copies of 3D toric codes. We apply the unfolding unitaries introduced in Section 2.4. To simplify the presentation of the boundary, we have opted for a different alignments from the one shown for the string operators. In this case, we consider the front and rear sides of the green (and purple) lattice to be the $Z$-boundaries, as opposed to the top and bottom sides as previously shown. The green sublattice is given by
@@ 3.2 The $X$-Boundary
Similar to the case with $Z$-boundaries, three independent types of membrane operators can terminate on this boundary and commute with every stabilizer. However, the supports of the membrane operators can differ depending on the specific cellulation. In our example, we display the configurations of these (logical) membrane operators in Fig. 4. We define the $X$-membrane operator on the $\mathbf{c}_{i}\mathbf{c}_{j}$-colored membrane as $\overline{X}_{\mathbf{c}_{i}\mathbf{c}_{j}}$ if it is also a logical operator, and define it as $\widetilde{X}_{\mathbf{c}_{i}\mathbf{c}_{j}}$ if it is not.
@@ 3.3 Other boundaries
3.3.1 The $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$-boundaries
@@
3.3.2 The $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$-boundaries
@@
where each two-body $Z$-type stabilizer is supported on an edge of a square on the boundary, and each four-body $X$-type stabilizer is supported on a full square on the boundary. We remark that the geometry of the boundary lattice is the same as that of the previously discussed $\{(\mathbf{c}_{i})_{\mathbf{z}},(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$-boundaries. However, by introducing the second stabilizer in Eq. (21), the second stabilizer in Eq. (19) is no longer preserved.
@@
3.3.3 The folded boundaries
Consider there are two copies of 3D toric codes, we can fold one copy of it, and attach the folded boundary to either the $e$-boundary, or the $m$-boundary of the second copy. We refer to this type of boundaries as the folded boundaries, implying they are created by folding toric codes. Since one copy of 3D toric code is folded along a codimension-1 (2D) submanifold in the bulk, the condensation set on the folded interface is $(e_{i}e_{j},m_{i}m_{j})$. Together with the boundary of the second copy, one can obtain two different types of boundaries, namely, the $(e_{i}e_{j},m_{i}m_{j},e_{k})$-boundary and the $(e_{i}e_{j},m_{i}m_{j},m_{k})$-boundary. According to our previous discussion, this boundary correspond to the 3D color-code boundary with condensations $\{(\mathbf{c}_{i})_{\mathbf{z}}(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{j}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{k})_{\mathbf{z}}\}$ and $\{(\mathbf{c}_{i})_{\mathbf{z}}(\mathbf{c}_{j})_{\mathbf{z}},(\mathbf{c}_{j}\mathbf{c}_{k})_{\mathbf{x}}(\mathbf{c}_{i}\mathbf{c}_{k})_{\mathbf{x}},(\mathbf{c}_{i}\mathbf{c}_{j})_{\mathbf{x}}\}$, respectively. For the simplicity of notations, we refer to the former boundary as the fold|e|e-boundary and the latter as the fold|m|m-boundary in the reminder of this paper, as these boundaries are composites of the fold of one copy of 3D toric code and the $e$-boundary or $m$-boundary of another copy, respectively.
Without loss of generality, we use explicity examples to show how to construct these two types of boundaries. For the fold|e|e-boundary, we consider the example when the condensation set is $\{y_{\mathbf{z}}g_{\mathbf{z}},pg_{\mathbf{x}}py_{\mathbf{x}},p_{\mathbf{z}}\}$. Just like the cases for 3D toric code, as we illustrated in Appendix B.2.3, we introduce ancilla qubits on the boundary to realize the condensation property [^4] Technically speaking, adding ancilla qubits is not necessary; one can multiply $X$-stabilizers to eliminate the need for ancilla qubits. However, ancilla qubits can serve as degrees of freedom that are purely defined on the boundary. By applying local operators on the boundary, typically single $X$ or $Z$ operators, one can observe how the condensations emerge from the boundary and become gapped in the bulk.. We introduce one ancilla qubit on each horizontal trancated yellow-green faces, as denoted by the white circles in the following figure.